OpenStudy (anonymous):

simplify 99 - 3(4 x 2^3)

OpenStudy (anonymous):

3

OpenStudy (angela210793):

it is 3 check the last post :)

jhonyy9 (jhonyy9):

99-3(4x8)=99-(3x32)=99-96=3

OpenStudy (anonymous):

ya 3!!!